OpenStudy (anonymous):

Find the inverse of each function. Is the inverse a function? To start, switch x and y. y= x/2

OpenStudy (ash2326):

Yeah if inverse exist for a function then it's also a function. For example here we have y= x/2 to find inverse of this replace x---->y y---->x so we have \[x=y/2\] now this can be written as \[y/2=x\] or \[y=2x\] so y=2x is the inverse function of y=x/2 to check if it's correct Let's put x=2 in y=x/2 we get y=2/2=1 now use this as x in the inverse function y=2x x=1 \[y=2\times 1\] y=2 so we began with 2 and got back 2 so y=2x is inverse of y=x/2

OpenStudy (shayaan_mustafa):

y=x/2 switch x and y so we get x=y/2 now solve for y y=2x (It is the required function.) Good Luck

OpenStudy (anonymous):

x=2y